### v1

# Advancing Spiking Neural Networks towards Multiscale Spatiotemporal Interaction Learning

Abstract
Recent advancements in neuroscience research have propelled the development of Spiking Neural Networks (SNNs), which not only have the potential to further advance neuroscience research but also serve as an energy-efficient alternative to Artificial Neural Networks (ANNs) due to their spike-driven characteristics. However, previous studies often neglected the multiscale information and its spatiotemporal correlation between event data, leading SNN models to approximate each frame of input events as static images. We hypothesize that this oversimplification significantly contributes to the performance gap between SNNs and traditional ANNs. To address this issue, we have designed a Spiking Multiscale Attention (SMA) module that captures multiscale spatiotemporal interaction information. Furthermore, we developed a regularization method named Attention ZoneOut (AZO), which utilizes spatiotemporal attention weights to reduce the model’s generalization error through pseudo-ensemble training. Our approach has achieved state-of-the-art results on mainstream neural morphology datasets. Additionally, we have reached a performance of 77.1% on the Imagenet-1K dataset using a 104-layer ResNet architecture enhanced with SMA and AZO. This achievement confirms the state-of-the-art performance of SNNs with non-transformer architectures and underscores the effectiveness of our method in bridging the performance gap between SNN models and traditional ANN models. All source code and models are available at https://github.com/Ym-Shan/Spiking_Multiscale_Attention_Arxiv.
1 Introduction
The biological brain has long served as a rich source of inspiration for the development of neural networks. By successfully mimicking the complex hierarchical structure of the visual cortex, Artificial Neural Networks (ANNs) have achieved numerous impressive accomplishments [1, 2]. However, the increasing energy consumption has become a major limitation hindering the further progress of ANNs. In comparison, Spiking Neural Networks (SNNs), which utilize binary spiking signals (0 for no spiking or 1 for spiking), possess inherent low-power characteristics. These low-power attributes of SNNs mainly arise from non-continuous activation [3] and spike-driven properties [4]. The emergence of neuromorphic chips such as Loihi [5] and TrueNorth [6] is expected to accelerate the widespread adoption of SNNs. The primary goal and ongoing challenge in SNNs research is to simultaneously draw inspiration from both high-performance deep learning models and the mechanisms of the biological brain in order to achieve brain-inspires intelligence.
In the initial stages of SNNs’ development, researchers encountered challenges in training algorithms. Drawing inspiration from synaptic plasticity in neuroscience and the backpropagation algorithm in deep learning, they proposed training methods such as STDP unsupervised learning [7], ANN2SNN [8, 9], and STBP [10]. Building upon this foundation, some researchers incorporated elements from deep learning, such as VGG [11] and ResNet [12, 13] architectures, into SNNs or optimized loss functions to enhance model performance. Additionally, motivated by insights from neuroscience, some researchers introduced attention mechanisms into SNNs [14, 15] or designed neuron simulation models that mimic the real properties of the biological brain, aiming to achieve more brain-inspires characteristics.
However, most prior studies have not delineated discrepancies in resolution and morphology among sample features within mainstream datasets. Consequently, the majority of these works have concentrated on devising comprehensive network architectures rather than considering multiscale structures in their designs. Furthermore, with the exception of TCJA [15], none of the works have taken into account the spatiotemporal correlations. Visualizing the spike dynamics of the bottleneck neurons in mainstream networks, as depicted in Fig. 1, we analyze their learning patterns. The results indicate that SNN models, which overlook multiscale information and spatiotemporal correlations, exhibit recognition patterns for event data akin to ANN networks recognizing static images. Moreover, SNN models relying solely on spatiotemporal correlations fail to alter this recognition pattern. We posit that this oversight has led to contemporary SNNs forfeiting some brain-inspires characteristics.
Therefore, we propose a Spiking Multiscale Attention (SMA) module to introduce multiscale representation learning into the SNN domain, while simultaneously leveraging spatiotemporal correlations to compute attention weights to address this issue. This method enhances the model’s ability to extract multiscale features while altering the learning pattern of SNNs, enabling SNN models to better balance local and global features. It is worth noting that we attribute the performance gap between SNNs and ANNs to the insufficient utilization of spatiotemporal correlation information by SNN models. To further mitigate this limitation, we devised an Attention Zoneout (AZO) regularization method. This method enhances model generalization by replacing information at spatiotemporal weak points of hidden units with information from previous time step. Unlike dropout, the AZO regularization method involves substitution rather than deletion, facilitating smoother propagation of gradient and state information through time. The primary contributions of this paper are summarized as follows:
We observe a neglect in previous works towards multiscale information and spatiotemporal correlation, and identify that this leads to the loss of brain-inspired features in SNN models, consequently leading to erroneous learning patterns.
We extend the SE module [16] to multiple scales and propose the SMA module based on it to introduce multiscale representation learning into SNNs, while utilizing spatiotemporal correlation information to compute attention weights for balancing local and global information, thereby steering SNN models towards a more brain-like learning pattern. To the best of our knowledge, this marks the first attempt to integrate multiscale representation learning into SNNs.
To further leverage spatiotemporal correlation information, we introduce an AZO regularization method leveraging spatiotemporal attention weights. This method employs the previous hidden unit value as noise to replace irrelevant information, thereby training a pseudo ensemble to enhance the robustness and generalization of the model.
We demonstrate the effectiveness of our proposed approach by achieving state-of-the-art accuracy on three mainstream datasets and Imagenet-1K. Through comprehensive visualization analysis, we provide evidence that the method of simultaneously leveraging multiscale information and spatiotemporal correlation can indeed induce SNNs to adopt a more brain-inspires learning pattern.
2 Related works
Attention Mechanism has emerged as a key factor in improving deep learning model performance, complementing depth, width, and cardinality. As traditional model structure design reaches a bottleneck, developing more effective attention mechanisms to focus on crucial features has become essential. In the SNN realm, Yao et al. introduced the TA module [17] as the inaugural attention mechanism emphasizing feature importance in the time dimension. Subsequently, the same team devised the MA module [14] to simultaneously highlight feature significance across time, spatial, and channel dimensions. From a different perspective, Zhu et al. proposed the tcja module [15], a portable and efficient solution to better concentrate on the characteristics of the time and channel dimensions, considering their interrelation. Shan et al. pioneered the integration of discrete spiking signals into attention-based decision-making processes [18]. Considering the need for lightweight design and the utilization of multiscale spatiotemporal correlation information, we developed the MSE module and the SMA module based on it, which transform the model’s learning pattern at minimal cost.
Multiscale Representation Learning plays a pivotal role in various computer vision applications, including image classification [19, 20, 21], object detection [22, 23, 24], image segmentation [25, 26], and image fusion [27, 28]. Given the diverse shapes and resolutions of natural objects, multiscale approaches enable the identification of relevant features [29]. With the advancement of deep learning, it has been observed that the depth of a network correlates positively with its feature representation capability [1, 30]. To enhance this capability, researchers have integrated multiscale representation learning into deep learning frameworks by designing multiscale convolution structures [31], incorporating pyramid structures [32, 33, 34], formulating multiscale loss functions [35, 36], and developing multiscale attention mechanisms [37, 38]. In this work, we broaden the scope of multiscale representation learning to SNNs by introducing a multiscale attention module.
Regularization reduces a model’s generalization error, typically dominated by variance, by balancing added bias against reduced variance. Data regularization methods like Cutout [39], Mixup [40], and CutMix [41] apply input image transformations to enhance model robustness. Structural regularization includes Dropout [42], adapted for specific tasks like Zoneout [43] for sequences and Dropblock [44] for segmentation, improving performance by selectively retaining unit values. Due to the approximate nature of dropout, researchers have explored optimal locations; Maxdropout [45] selects based on neuron activation, while Autodropout [46] chooses based on training insights. Our AZO regularization method utilizes SMA attention weights for selection, replacing irrelevant information with the hidden unit values from previous timesteps as noise, thereby training a pseudo-ensemble to enhance robustness and generalizability.
3 Method
3.1 Leaky Integrate-and-Fire Neuron Model
The Leaky Integrate-and-Fire (LIF) neuron [3] stands as one of the predominant neuron models within SNNs, esteemed for its balanced performance and adherence to biological principles, is uniformly adopted in this work. Within neural networks, neurons serve as fundamental computational units. Upon receiving transmitted spiking signals, LIF neurons initiate an integration process. Upon reaching a threshold membrane potential, neurons emit pulses and reset their membrane potentials. This dynamic process is encapsulated by the subthreshold dynamics model [4]
where represents a time constant, denotes the membrane potential of the postsynaptic neuron, and signifies the input gathered from presynaptic neurons. Additionally, denotes the reset potential, which is established subsequent to the activation of the output spiking.To facilitate training and description, we adopt the displayed iteration version of the subthreshold dynamics model [47].
at each layer indexed by and time step , the membrane potential of a neuron is denoted as . The parameter signifies a time constant, and represents a binary spiking tensor. denotes the neuron’s input, while represents the Heaviside step function. symbolizes the hidden state, refers to the reset potential of the neuron following a spike, and indicates the discharge threshold of the neuron.
3.2 Spiking Multiscale Attention for SNNs
The overall structure of the SMA module is illustrated in Fig. 2. Its primary objective is to enhance the model’s ability to extract target features with diverse resolutions, shapes, and states by utilizing spatiotemporal correlation information. This approach helps balance the significance of global and local features, thereby transforming the model’s learning pattern. The temporal and spatial information inherent in event data can often be intertwined, posing additional challenges in extracting multiscale features. Hence, instead of directly conducting multiscale feature extraction within the encoder, we opt to perform multiscale coding beforehand to enhance feature representation. The encoder’s architecture is delineated in the right portion of Fig. 2. Given an input event sequence , the encoder of the SMA module across scales can be represented as
where represents the multiscale coding process and denotes the function. However, when analyzing the mechanism of SMA, we substitute LIF neurons for (we have conducted ample experiments in Sec. C.3 of the appendix demonstrating that LIF does not notably compromise model accuracy compared to ). indicates that the 2D convolution operation is performed on using the convolution kernel , and represents the multiscale coding result of the t-th frame of the input event.
The decoder structure is illustrated in the lower-left portion of Fig. 2. It is responsible for processing the multiscale data , enriched with characteristics output from the encoder. Specifically, it calculates the attention weights for the temporal and channel dimensions as follows:
where and respectively describe the roles of the MSE (T-MSE) module in the time dimension and the MSE (C-MSE) module in the channel dimension. and represent the Squeeze convolution kernel () in the T-MSE module and C-MSE module, respectively. and represent the Excitation convolution kernel () in the n-th scale, respectively. represents the operation.
To compute attention weights in the time dimension, we adhere to the conventional methodology: initially conducting global average pooling on the input data, followed by computing attention weights to derive weights sized . To comprehensively leverage the interplay between multiscale information and spatiotemporal interactions, attention weights are computed for all input channels at each timestep. It is noteworthy that upon completing the calculation, operations are performed on the weights within the time and channel dimensions correspondingly, in the scale dimension, to mitigate the adverse effects of large (small) values on the model. Consequently, are obtained.Finally, apply the attention weights to the multiscale encoding result of the input event stream and aggregate them along the scale dimension:
3.3 Attention Zoneout for SMA
Zhu et al. were the first to explore the correlation between temporal and spatial information in event data within the domain of SNNs [15]. However, most mainstream SNN methods overlook this spatiotemporal correlation information. To address this, we propose a regularization method called Attention Zoneout (AZO), which utilizes temporal and channel attention weights, as described in Algorithm 1, to leverage these valuable pieces of information.
Similar to Zoneout [43], AZO trains pseudo-ensembles by adding noise to hidden units, thereby enhancing generalization ability. The key difference is that the location of noise in AZO is determined by attention weights rather than being random, which helps mitigate the impact of irrelevant features while introducing noise, thus facilitating the network’s convergence to the global optimum. Additionally, while Zoneout directly adds noise in the spatial dimension, AZO applies noise in the channel dimension because, at the time of AZO operation, all features have already been aggregated onto the channel dimension, resulting in a spatial dimension size of . Consequently, each operation applied to the channel dimension is equivalently projected onto the spatial dimension. In Sec. C.2 of the appendix, we present an ablation study to elucidate the selection process of the hyperparameters and in AZO and provide additional experiments to validate its efficacy.
4 Experiments
Even with the use of spiking coding, the spatiotemporal interaction information in static image datasets remains limited. Therefore, we only evaluated the classification performance of our proposed SMA-SNN and SMA-AZO-SNN architectures on three prominent neuromorphic datasets (DVS128 Gesture [48], CIFAR10-DVS [49], and N-Caltech101 [50]) and the ImageNet-1K dataset [51], as elaborated in Sec. 4.2. Preceding this assessment, we provided a comprehensive exposition of our attention position determination and scale quantity selection process in Sec. 4.1, a crucial aspect we regard as essential for a generic attention module. All network structures and hyperparameter settings utilized in the experiment are detailed in Sec. B of appendix.
4.1 Ablation study
Building upon the foundation established by prior research, it has been observed that the classification performance of SNN models on the DVS128 Gesture dataset [48] can be effectively extrapolated to other prominent datasets [17, 14, 15, 12]. Consequently, our ablation studies will be systematically conducted utilizing the DVS128 Gesture dataset as the primary benchmark.
SMA Position. Incorporating the plug-and-play attention module into mainstream network architectures necessitates careful consideration of its insertion points and quantity. Initially, we address the unique characteristics of the encoding module (the first module of SNNs) and categorize the potential roles of SMA within the network as follows: T1, absence of SMA; T2, exclusive addition of SMA to the encoding module; T3, integration of SMA throughout the network except for the encoding module; T4, comprehensive inclusion of SMA across the entire network. As depicted in Fig. 3, T3 consistently outperforms the other configurations across nearly all simulation timesteps, particularly at the prevalent timestep of 16.
Building upon the aforementioned findings, we further subdivide the potential optimal positions for SMA within the network as follows: S1, incorporating SMA solely in odd-numbered blocks excluding the encoding module; S2, integrating SMA exclusively in even-numbered blocks; S3, embedding SMA only in the initial half of the network, but excluding the coding module; S4, inserting SMA solely in the latter half of the network. As illustrated in Fig. 3, the S3 configuration demonstrates superior performance compared to other positions across all simulated timesteps. Additionally, we investigated the influence of SMA within the conventional SNN module (Conv-BN-Neuron). The potential insertion points for SMA are: L1, immediately following the convolution layer; L2, following the BN layer; L3, after the neuron layer. The experimental findings demonstrate that within the module, the placement of the SMA module at position L1 exhibits a slight advantage over its placement at L3, while positioning it at L2 results in the network’s failure to converge. This phenomenon is hypothesized to stem from the SMA module’s inability to accurately extract significance from normalized data, consequently leading to the disappearance of the network’s gradient.
| Scale | Acc(%) | Inference Overhead(s) |
| No SMA | 96.52 | 18.98 |
| 2 | 96.52(0.34) | 21.99(0.98) |
| 3 | 97.21(0.69) | 23.02(0.07) |
| 4 | 97.91(0.34) | 27.44(0.24) |
| 5 | 97.73(0.17) | 33.72(0.2) |
| 6 | 97.56(0.34) | 42.43(0.06) |
Scale Quantities. Paying attention to various receptive fields of differing sizes is crucial for SMA. The results in Tab. 1 demonstrate that the SMA module, integrating downsampling and attention mechanisms utilizing filters of sizes 1, 3, 5, and 7 across four distinct scales, yields the most pronounced and consistent effects. Furthermore, this configuration introduces only marginal additional reasoning time, thereby facilitating efficient processing for individual events without a significant increase in computational load.
4.2 Comparison with the State-of-the-Art
| Work | Architecture | Timestep | Top-1 Acc.(%) |
| SEW ResNet [12] | SEW-ResNet-34 | 4 | 67.04 |
| SEW-ResNet-50 | 4 | 67.78 | |
| SEW-ResNet-101 | 4 | 68.76 | |
| SEW-ResNet-152 | 4 | 69.26 | |
| MS ResNet [13] | MS-ResNet-18 | 6 | 63.10 |
| MS-ResNet-34 | 6 | 69.42 | |
| MS-ResNet-104 | 5 | 74.21 | |
| MS-ResNet-104* | 5 | 76.02 | |
| MA-ResNet [14] | MA-MS-ResNet-18 | 1 | 63.97 |
| MA-MS-ResNet-34 | 1 | 69.15 | |
| MA-MS-ResNet-104* | 4 | 77.08 | |
| Spiking ResNet [52] | ResNet-50 | 350 | 72.75 |
| Hybrid training [53] | ResNet-34 | 250 | 61.48 |
| TET [54] | SEW-ResNet-34 | 4 | 68.00 |
| tdBN [55] | Spiking-ResNet-34 | 6 | 63.72 |
| Spike-Norm [11] | ResNet34 | 2000 | 65.47 |
| QCFS [56] | VGG-16 | 64 | 72.85 |
| Fast-SNN [57] | VGG-16 | 7 | 72.95 |
| ResNet [13] | Res-CNN-104 | 1 | 76.87 |
| Spikformer [58] | Spiking Transformer-10-512 | 4 | 73.68 |
| Spiking Transformer-8-768* | 4 | 74.81 | |
| Spike-driven Transformer [59] | Spiking Transformer-10-512 | 4 | 74.66 |
| Spiking Transformer-8-768 | 4 | 77.07 | |
| Meta-SpikeFormer [60] | Meta-SpikeFormer | 4 | 77.20 |
| Meta-SpikeFormer* | 4 | 80.00 | |
| SMA-ResNet(Ours) | SMA-MS-ResNet-18 | 6 | 68.46 |
| SMA-MS-ResNet-34 | 6 | 70.19 | |
| SMA-AZO-MS-ResNet-104* | 5 | 77.05 |
Experimental results on Imagenet-1K. We adopted the same approach as Yao et al [14], simply integrating the SMA module into MS-ResNet [13] to evaluate its performance. Tab. 2 presents the experimental results. The integration of SMA alone significantly improved model accuracy, evidenced by enhancements of 5.36% in the 18-layer ResNet and 0.77% in the 34-layer ResNet configurations. Furthermore, the combined application of SMA and AZO resulted in an additional accuracy increase of 1.03% in the 104-layer structure. Comparison with previous studies indicates that our approach achieves state-of-the-art accuracy in traditional convolutional-based SNNs and is also competitive with some models based on the Transformer architecture.
| Work | Spike-driven | DVS128 Gesture | CIFAR10-DVS | N-Caltech101 | |||||
| T | Acc | T | Acc | T | Acc | ||||
| PLIF [61] | ✓ | 20 | 97.6 | 20 | 74.8 | - | - | ||
| Spikformer [58] | ✗ | 16 | 98.3 | 16 | 80.9 | - | - | ||
| tdBN [55] | ✗ | 40 | 96.9 | 10 | 67.8 | - | - | ||
| SEW-ResNet [12] | ✗ | 16 | 97.9 | 16 | 74.4 | - | - | ||
| TA-SNN [17] | ✗ | 60 | 98.6 | 10 | 72.0 | - | - | ||
| HATS [62] | N/A | - | - | N/A | 52.4 | N/A | 64.2 | ||
| DART [63] | N/A | - | - | N/A | 65.8 | N/A | 66.8 | ||
| SALT [64] | ✗ | - | - | 20 | 67.1 | 20 | 55.0 | ||
| TCJA-SNN [15] | ✗ | 20 | 99.0 | 10 | 80.7 | 14 | 78.5 | ||
✓ | 16 | 99.3 | 16 | 80.0 | - | - | ||
| SMA-VGG(Ours) | ✓ | 16 | 98.3 | 10 | 83.1 | 14 | 83.7 | ||
| SMA-AZO-VGG(Ours) | ✓ | 16 | 98.6 | 10 | 84.0 | 14 | 84.6 | ||
Experimental results on mainstream neuromorphic datasets. As Tab. 3 illustrates, the integration of SMA into the VGG network allows for SOTA performance on the CIFAR10-DVS and N-Caltech101 datasets. The accuracy is further and steadily improved through the application of the AZO regularization method. Our approach also resulted in a significant 6.1% increase in accuracy on the N-Caltech101 dataset. Moreover, on the DVS128 Gesture dataset, our method achieved equivalent results to those of non-transformer architectures with the same timesteps.
4.3 Study of Learning Patterns
SMA Changes SNNs Learning Patterns. Observations of the attention heatmaps depicted in Fig. 1 and 4(b) reveal that traditional SNN decision-making approaches treat each frame of event data as a static image. This strategy leads to an overemphasis on global features and a neglect of local features, thereby impeding the utilization of temporal information. However, as depicted in Fig. 4(c), the incorporation of the SMA module balances the importance of local and global features by integrating multiscale and spatiotemporal correlation information. This shift in the model’s decision-making mode towards recognizing the relative positions and dynamic information of crucial joints enhances the utilization of spatiotemporal correlation information between event frames.
The N-Caltech101 dataset has more complex features and more noise. A closer examination of Fig. 4 reveals that SNN models lacking SMA tend to allocate unnecessary attention to the background while neglecting local features. We identify this as the primary reason for the underperformance of traditional SNN models on the N-Caltech101 dataset. Upon integrating SMA, the SMA-SNN model effectively mitigates the issue of background overemphasis observed in traditional SNNs, redirecting attention more precisely towards relevant local features. This enhancement is believed to be the primary factor contributing to the improved performance of the SMA-SNN model on the N-Caltech101 dataset. Additionally, as depicted in Fig. 4 right, SMA significantly reduces the SFR of the SNN model by minimizing unnecessary attention to the background, thereby promising greater energy savings for event-driven SNNs.
4.4 Study on the correlation between SMA and multiscale information in samples
Focus on multiscale features. The filters of sizes 1, 3, 5, and 7 in Fig. 5 are employed for downsampling across the four scales. Research conducted on this figure suggests that in the decision-making process, samples characterized by smaller key features (such as fingers and wrists) exhibit a preference for filters of smaller size, whereas samples with larger key features (such as arms and trajectories) show a predilection for filters of larger size. This observation underscores the intrinsic capability of the SMA module to enhance model performance in datasets with a diverse range of feature sizes.
5 Conclusion
In our observations of previous work, we found that SNN models that ignore multiscale information and spatiotemporal correlation information exhibit recognition patterns for event data similar to ANN models recognizing static images. Additionally, visualizations of the heatmaps from previous work indicate that solely relying on spatiotemporal correlation information cannot fundamentally alter the recognition patterns of SNNs. Therefore, we introduced multiscale spatiotemporal interaction learning into SNNs by designing the Spiking Multiscale Attention (SMA) module to alter the learning patterns of SNNs. Experiments demonstrate that SMA effectively utilizes both multiscale information and spatiotemporal correlation information to balance the importance of global and local features in the samples, thereby changing the learning patterns of SNNs. We further utilize spatiotemporal correlation information by proposing an AZO regularization method. This method replaces the information at spatiotemporal weak points with the information from the corresponding spatial locations at the previous time step to train pseudo-ensembles, effectively reducing the model’s generalization error. SMA and AZO achieve state-of-the-art accuracy on CIFAR10-DVS (84.0%) and N-Caltech101 (84.6%) datasets, while also attaining non-Transformer architecture state-of-the-art accuracy on the ImageNet-1k dataset (77.1%), underscoring their efficacy. Therefore, our work highlights the importance of multiscale information and spatiotemporal correlation information in SNNs, advancing SNNs one step closer to the decision-making processes of the human brain.
APPENDIX
Appendix A Detail of Datasets
A.1 Datasets
This article uses three mainstream neural morphology datasets(DVS128 Gesture, CIFAR10-DVS and N-Caltech 101) and the Imagenet-1K dataset.
DVS128 Gesture is an event-based dataset comprising a sequence of 11 gestures. It consists of 1176 samples in the training set and 288 samples in the test set. The training set was captured by 23 subjects, while the test set was captured by 6 subjects, each under three different lighting conditions. The samples are sized at 128 128 pixels.
CIFAR10-DVS is a conversion from the CIFAR10 dataset, consisting of ten categories totaling 10000 images. The size of each frame has also been expanded to 128 128 pixels. Due to the additional undesired artifacts included in the neuromorphic datasets obtained based on the conversion method, therefore, how to distinguish the categories of targets in complex backgrounds with additional errors is a challenging recognition task.
N-Caltech 101 presents a notable challenge, arising from the transformation and slight adaptation of the Caltech dataset to avoid confusion. Comprising 8,709 samples, it is segmented into 100 distinct object classes alongside 1 background class. Emarkably, its image dimensions (180 240) surpass those of numerous other neural morphology datasets.
Imagenet-1K is the most renowned and extensively utilized static image classification dataset. It comprises 1.28 million training samples and 50,000 testing samples, distributed across 1,000 different categories. Each image in the dataset has a resolution of 224x224 pixels.
For the partitioning of datasets like CIFAR10-DVS and N-Caltech101, which necessitate division, we adopt methodologies from prior studies and partition them in a 9:1 ratio. We will furnish the necessary .py scripts for facilitating the partitioning process.
A.2 Neuromorphic Datasets Preprocessing
We use the integrating approach to convert event stream for frame data, which is commonly used in SNNs, to preprocess neuromorphic datasets. The coordinate of event can be describe as
where and event’s coordinate, represents the event. In order to deduce computational consumption, we group events into slices, Where is the network’s time simulation step. The process can be describe as
where is the floor operation, and is an indicator function and it equals 1 only when .
A.3 Data augmentation
| Dataset | Data augmentation | ||||
| CIFAR10-DVS |
| Imagenet-1K |
For the DVS128 Gesture and N-Caltech101 datasets, we adhered to our previous experience and refrained from applying data augmentation. To address the significant overfitting issue observed on the CIFAR10-DVS and Imagenet-1K datasets, we implemented data augmentation strategies outlined in Tab. 4.
| Block | Output Size | SMA-ResNet-18 | SMA-ResNet-34 | SMA-ResNet-104 |
| 1 | 112 112 | 7 7, 64, stride=2 | ||
| 2 | 56 56 | SMA(CR=4,TR=1) | SMA(CR=4,TR=1) | SMA(CR=4,TR=1) |
| 3 | 28 28 | SMA(CR=4,TR=1) | SMA(CR=4,TR=1) | SMA(CR=4,TR=1) |
| 4 | 14 14 | SMA(CR=4,TR=1) | SMA(CR=4,TR=1) | SMA(CR=4,TR=1) |
| 5 | 7 7 | SMA(CR=4,TR=1) | SMA(CR=4,TR=1) | SMA(CR=4,TR=1) |
| FC | 1 1 | AveragePool, FC(1000) | ||
| Dropout | 1 1 | Null | Dropout(0.2) | |
Appendix B Training details
B.1 Network structure
| Block | SMA-VGG | Output Size | ||
| DVS128 Gesture | CIFAR10-DVS | N-Caltech101 | ||
| 1 | 3 3,64 | 128 128 | 128 128 | 180 240 |
| MaxPool(2,2,0,1) | 64 64 | 64 64 | 90 120 | |
| 2 | 3 3,128 | 32 32 | 32 32 | 45 60 |
| SMA(CR=4,TR=4) | ||||
| MaxPool(2,2,0,1) | ||||
| 3 | 3 3,256 | 16 16 | 16 16 | 22 30 |
| SMA(CR=4,TR=4) | ||||
| MaxPool(2,2,0,1) | ||||
| 4 | 3 3,512 | 8 8 | 8 8 | 11 15 |
| SMA(CR=4,TR=4) | ||||
| MaxPool(2,2,0,1) | ||||
| 5 | 3 3,512 | 4 4 | 4 4 | 5 7 |
| SMA(CR=4,TR=4) | ||||
| MaxPool(2,2,0,1) | ||||
| FC-1 | AveragePool | 1 1 | 1 1 | 1 1 |
| FC(2048) | ||||
| Dropout(0.5) | ||||
| FC-2 | FC(1024) | 1 1 | 1 1 | 1 1 |
| Dropout(0.5) | ||||
| FC-3 | FC(11/10/101) | 1 1 | 1 1 | 1 1 |
In the experiment detailed in Chap. 4 of the primary text, we employed MS-ResNet networks with varying numbers of layers on the ImageNet-1K dataset, as illustrated in Tab. 5. Additionally, we utilized an 8-layer VGG architecture on the mainstream neural morphology dataset, as depicted in Tab. 6. It is worth noting that in both the SMA-ResNet and SMA-VGG architectures, our SMA module closely follows the convolutional layer, and an SMA module is added after each residual connection. This suggests that the actual configuration of the second module in SMA-ResNet18 consists of Conv-SMA-Conv-Conv-SMA-Conv, while the actual structure of the second module of SMA-VGG is Conv-SMA-MaxPool. For clarity, we omitted LIF neurons and BatchNorm in Tab. 5 and 6. The MS residual block structure in Tab. 5 is detailed in Fig. 6, and each convolution operation in Tab. 6 can be interpreted as LIF Neuron-Conv-BN.
B.2 Training strategy
| Model | Dataset | Name | Value |
| SMA-VGG | DVS128 Gesture | lr | 1e-4 |
| T | 16 | ||
| batch_size | 9 | ||
| train_epoch | 200 | ||
| loss_function | MSE_loss | ||
| optimizer | AdamW(momentum_decay=1e-3) | ||
| CIFAR10-DVS | lr | 1e-3 | |
| T | 10 | ||
| batch_size | 24 | ||
| train_epoch | 200 | ||
| loss_function | MSE_loss | ||
| optimizer | Adam | ||
| N-Caltech101 | lr | 1e-3 | |
| T | 14 | ||
| batch_size | 6 | ||
| train_epoch | 300 | ||
| loss_function | TET_loss | ||
| optimizer | NAdam(momentum_decay=1e-3) | ||
| SMA-ResNet18/34 | Imagenet-1K | lr | 0.1 |
| T | 6 | ||
| batch_size | 384/256 | ||
| train_epoch | 125 | ||
| loss_function | Label_smoothing(0.1) | ||
| optimizer |
| SMA-ResNet104 | Imagenet-1K | lr | 0.1(1e-3) |
| T | 5 | ||
| batch_size | 224 | ||
| train_epoch | 125(50) | ||
| loss_function | Label_smoothing(0.1) | ||
| optimizer |
| Datasets | Parameter | Value |
| Neuromorphic Datasets | 1.0 | |
| 0.0 | ||
| 2.0 | ||
| Surrogate Gradient Function | ATan() | |
| Imagenet-1K | 0.5 | |
| 0.0 | ||
| 4.0 | ||
| Surrogate Gradient Function |
The experiments conducted on the neuromorphic datasets in this article were all carried out using 4 3090GPUs, while experiments on Imagenet-1K were performed using 6/8/10 3090D GPUs depending on the ResNet layers. Tab. 7 presents the hyperparameters necessary for training each dataset, while Tab. 8 delineates the membrane constants of LIF neurons for both the neuromorphic dataset and Imagenet-1K dataset.
B.3 Equipment details
| No | CPU | GPU | Memory | CUDA | Pytorch | OS |
| 1 | Gold 6133 2 | RTX 3090 4 | 128GB | 12.1 | 2.0.1 | Ubuntu 18.04 |
| 2 | Gold 6430 | RTX 4090 8 | 960GB | 12.1 | 2.1.0 | Ubuntu 22.04 |
| 3 | AMD EPYC 9754 | RTX 4090D 10 | 600GB | 12.1 | 2.1.0 | Ubuntu 22.04 |
Appendix C Hyperparameter selection
For general attention modules, particularly those based on SE modules, the hyperparameters within the modules are crucial, as they frequently have the potential to influence the performance of attention mechanisms to a certain extent.
C.1 SMA hyperparameter selection
The SMA module features only two hyperparameters: the channel reduction ratio and the time reduction channel. As in Chap. 4 of the primary text, we undertook experiments using the DVS128 Gesture dataset to ascertain the optimal reduction ratios in both the channel and time dimensions. Initially, we integrated SMA solely along the channel dimension into the network for experimentation, with the results depicted in Fig. 7(a). Notably, setting the reduction ratio in the channel dimension to 4 yielded the most favorable outcomes. Building upon this finding, we further investigated the optimal reduction ratio in the time dimension, as illustrated in Fig. 7(b), where the optimal setting remained at 4.
The insights from Fig. 7 highlight the critical role of hyperparameter selection within the attention module, as an incorrect choice can detrimentally impact model performance. Upon close examination of the findings in Fig. 7, it becomes evident that excessive compression may blur the importance of information within the attention module, while insufficient compression may prevent the model from achieving the correct balance of importance.
C.2 AZO hyperparameter selection
The AZO regularization method utilizes two hyperparameters: the Replacement Time Ratio(RTR) and the Replacement Channel Ratio(RCR). Given the short simulation timesteps in our experiment, the influence of the timestep on the AZO effect appears to be considerably less significant than the number of channels, a notion supported by Fig. 9. Given the diverse nature of the datasets involved—CIFAR10-DVS and N-Caltech101, which are converted datasets, and DVS128 Gesture, which represents natural neural morphology—we investigated the optimal RCR for the DVS128 Gesture and CIFAR10-DVS datasets, setting the RTR at 4, as illustrated in Fig. 8. Following this, we further explored the optimal RTR for the DVS128 Gesture dataset, as depicted in Fig. 9.
| AZO hyperparameters | DVS128 Gesture | CIFAR10-DVS | N-Caltech101 |
| Best RCR | 10/12 | 4 | 24 |
| Best RTR | 4 | 4 | 4 |
We posit that the distinct data formats and acquisition methods necessitate uniquely suitable hyperparameter settings for each dataset when using the AZO regularization method. Due to the extensive data volume, we refrained from conducting an exhaustive exploration of optimal hyperparameters on the N-Caltech101 dataset, opting instead to provide a set of parameters that demonstrated improved performance, as detailed in Tab. 10.
C.3 Comparison of accuracy between LIF based and ReLU based SMA
| Type | DVS128 Gesture | CIFAR10-DVS | N-Caltech101 | |||
| SMA-C | SMA | SMA-C | SMA | SMA-C | SMA | |
| ReLU | 97.9 | 98.3 | 82.3 | 83.1 | 82.7 | 83.7 |
| LIF Neuron | 97.9 | 98.3 | 81.8 | 82.6 | 82.5 | 82.9 |
In the Discussion section of the primary text, we conducted extensive experiments utilizing the LIF version of the SMA module. The discrete nature of signals in SNNs facilitated easy visualization and analysis of attention modules based on spike count or firing rate—a method widely adopted in numerous studies. As illustrated in Tab.11, our experiments showcased that while the ReLU version of the SMA module exhibits slightly superior performance, the disparity in performance between it and the LIF version is minimal. Notably, even on the DVS128 Gesture dataset, both versions yielded identical effects. Hence, it is reasonable to posit that the LIF version and the ReLU version of the SMA module share the same attention mechanism, affirming our discussion in the primary text.
Appendix D Further discussion
D.1 Complexity
For an efficient plug-and-play attention module, it’s crucial to conduct a detailed analysis of both time and space complexities. In this chapter, we specifically concentrate on the process involved in computing attention weights, setting aside the Multiscale Coding process. The latter is primarily composed of straightforward convolution and addition operations, which are relatively less complex. For Channel Multiscale Attention and Time Multiscale Attention in SMA. In section 3.2 of the primary text, we describe them as follows:
| Dataset | Model | Inference time(s) | |
| No SMA | With SMA | ||
| DVS128 Gesture | VGG | 2.10 | 2.97 |
| CIFAR10-DVS | VGG | 3.26 | 4.79 |
| N-Caltech101 | VGG | 7.29 | 10.91 |
| Imagenet-1K | ResNet18 | 14.78 | 25.41 |
| ResNet34 | 19.34 | 29.87 | |
| ResNet104 | 39.99 | 52.25 | |
Time complexity. In analyzing the Time Multiscale Attention within SMA for multiscale encoded data sized , the Avgpool operation has a time complexity of . Post-Avgpool, the data size is reduced to . For this input data, the T-MSE Block initiates with a convolution operation, which has a time complexity of . It then proceeds to filter out negative values using the ReLU function, also with a time complexity of . Following this, another convolution is performed at various scales, with its own time complexity of . For processing the output results from the T-MSE Block, a SoftMax operation is necessary, adding a further time complexity of . Consequently, the total time complexity for the Time Multiscale Attention module sums up to .
Channel Multiscale Attention operates similarly to Time Multiscale Attention but without the Avgpool process. Instead, attention weights are computed at each timestep through the C-MSE Block. Consequently, the overall time complexity for Channel Multiscale Attention is . Given that and often represent smaller values, SMA only slightly increases inference time, and its additional inference time is independent of the depth of the network, as demonstrated in Tab. 12.
Spatial complexity. When calculating attention weights, the sole additional computational units that introduce parameters are the 1x1 convolutional layers in the T-MSE and C-MSE blocks. The spatial complexity for T-MSE is , and for C-MSE, it is . The additional parameters required by the convolution operation in multiscale encoders are unavoidable. However, the attention weight calculation (decoder) introduces only a minimal number of parameters.
D.2 Optimizing the implementation of AZO
| Version | Training time(s) | ||
| Dvs128 Gesture | CIFAR10-DVS | N-Caltech101 | |
| Explicit Loop | 55.29 | 802.54 | 755.07 |
| Tensor Parallelism | 4.35 | 68.61 | 44.04 |
Clearly, the efficiency of implementing AZO regularization in Python using nested display loops is quite low. Thus, we propose a method for AZO implementation that leverages the tensor parallelism of Python and Numpy. As illustrated in Tab. 13, we compared the time required for a single training iteration between the parallel and the loop-based versions of AZO. We found that the loop version offers a significant efficiency advantage, establishing a solid foundation for the widespread adoption of AZO-like regularization algorithms.
D.3 AZO effect
As elucidated in Section 3.3 of the primary text, the incorporation of the AZO algorithm is posited to enhance the generalizability of the model. This postulate is grounded in the algorithm’s methodology, which perturbs the hidden unit activations with noise during the training of pseudo-sets. To empirically substantiate this hypothesis, we employed the N-Caltech101 dataset—a dataset of considerable complexity, meticulously selected from among a consortium of smaller datasets. Fig. 10 provides a graphical representation of the empirical results, contrasting the model’s performance on training and testing accuracies both pre and post the integration of the AZO algorithm. The results unambiguously indicate that the AZO algorithm materially amplifies the generalization capabilities of the model.
D.4 The effect of SMA and AZO on ResNet Structure
| Work | DVS128 Gesture | CIFAR10-DVS |
| SMA-ResNet18 | +0.69% | +3.46% |
| SMA-AZO-ResNet18 | +1.04% | +4.96% |
In the main text, we exclusively showcased the efficacy of SMA within the ResNet architecture on extensive datasets such as Imagenet-1K. However, in this section, we extend our analysis to demonstrate its impact on smaller-scale datasets. Tab. 14 illustrates that SMA yields notable improvements within the ResNet framework even for datasets of reduced scale.
NeurIPS Paper Checklist
1.
Claims
Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
Answer: [Yes]
Justification: Our abstract accurately captures the motivation, methods, and results of our study. It has been validated through experiments using VGG and ResNet architectures across four datasets, including the ImageNet dataset, demonstrating strong generalization capabilities.
Guidelines:
The answer NA means that the abstract and introduction do not include the claims made in the paper.
The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
2.
Limitations
Question: Does the paper discuss the limitations of the work performed by the authors?
Answer: [Yes]
Justification: In the appendix, we have detailed the process of exploring hyperparameters for the SMA. The experiments clearly show that the performance of SMA modules is influenced by hyperparameters. Furthermore, the AZO regularization method depends on the attention weights of the SMA module.
Guidelines:
The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
The authors are encouraged to create a separate "Limitations" section in their paper.
The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
3.
Theory Assumptions and Proofs
Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
Answer: [N/A]
Justification: Our work does not belong to the field of deep learning theory research.
Guidelines:
The answer NA means that the paper does not include theoretical results.
All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
All assumptions should be clearly stated or referenced in the statement of any theorems.
The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
Theorems and Lemmas that the proof relies upon should be properly referenced.
4.
Experimental Result Reproducibility
Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
Answer: [Yes]
Justification: In the main text, we have extensively described the complete structure of our proposed method using visual and mathematical representations. Additionally, we have listed all hyperparameters for the experiments in the appendix. Moreover, we have made our source code publicly available.
Guidelines:
The answer NA means that the paper does not include experiments.
If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
5.
Open access to data and code
Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
Answer: [Yes]
Justification: Our experiments were conducted using publicly available datasets. Additionally, we have made all source code and weight files publicly accessible.
Guidelines:
The answer NA means that paper does not include experiments requiring code.
Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
6.
Experimental Setting/Details
Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
Answer: [Yes]
Justification: All hyperparameters are detailed in the appendix.
Guidelines:
The answer NA means that the paper does not include experiments.
The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
The full details can be provided either with the code, in appendix, or as supplemental material.
7.
Experiment Statistical Significance
Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
Answer: [Yes]
Justification: We conducted numerous experiments during the hyperparameter exploration process for the SMA and AZO, recognizing the significance of these for a universal module. The results in Tab. 1 of the main text and Tab. 8 in the appendix are averages from ten experiments using non-fixed random seeds. Similarly, the data shown in Fig. 5 of the appendix were derived from five experiments with non-fixed random seeds, including error bars to depict variability.Due to constraints on experimental costs and computational resources, we limited our testing to one experiment each on the ImageNet dataset and mainstream neural architecture datasets. However, we have carefully preserved the model weights for future reference and use.
Guidelines:
The answer NA means that the paper does not include experiments.
The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
The assumptions made should be given (e.g., Normally distributed errors).
It should be clear whether the error bar is the standard deviation or the standard error of the mean.
It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
8.
Experiments Compute Resources
Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
Answer: [Yes]
Justification: In the main text, we detailed the specific configurations of the three servers used, as well as the service providers for two of the cloud servers. All experiments and preliminary work described in this article, including the SMAL module that focuses solely on multiscale interactions without considering spatiotemporal dynamics, have been made open-source.
Guidelines:
The answer NA means that the paper does not include experiments.
The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
9.
Code Of Ethics
Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
Answer: [Yes]
Justification: Our research, while inspired by the biological brain, does not involve real organisms and contains no generative content, ensuring compliance with ethical standards.
Guidelines:
The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
10.
Broader Impacts
Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
Answer: [No]
Justification: We believe that our proposed method will drive the deployment of SNNs on brain like chips and the study of high-performance SNN models, which we consider to have a social impact. However, we did not discuss this in the main text because brain like chips are still in the research stage and their application fields are more inclined towards professional fields. We believe that our research has limited social impact beyond the academic field, so we did not discuss it.
Guidelines:
The answer NA means that there is no societal impact of the work performed.
If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
11.
Safeguards
Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
Answer: [N/A]
Justification: Our research is purely academic and does not involve generative studies; therefore, it carries no associated risks.
Guidelines:
The answer NA means that the paper poses no such risks.
Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
12.
Licenses for existing assets
Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
Answer: [Yes]
Justification: The four datasets we use are publicly available for academic research, and we have also cited them.
Guidelines:
The answer NA means that the paper does not use existing assets.
The authors should cite the original paper that produced the code package or dataset.
The authors should state which version of the asset is used and, if possible, include a URL.
The name of the license (e.g., CC-BY 4.0) should be included for each asset.
For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
13.
New Assets
Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
Answer: [Yes]
Justification: Yes, we have made our proposed model publicly available, and all related documentation has also been released.
Guidelines:
The answer NA means that the paper does not release new assets.
Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
The paper should discuss whether and how consent was obtained from people whose asset is used.
At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
14.
Crowdsourcing and Research with Human Subjects
Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
Answer: [N/A]
Justification: This study did not involve any subjects.
Guidelines:
The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
15.
Institutional Review Board (IRB) Approvals or Equivalent for Research with Human Subjects
Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
Answer: [N/A]
Justification: This study did not involve any subjects.
Guidelines:
The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.

### v2

@@
 14 illustrates that SMA yields notable improvements within the ResNet framework even for datasets of reduced scale.
-NeurIPS Paper Checklist
-1.
-Claims
-Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
-Answer: [Yes]
-Justification: Our abstract accurately captures the motivation, methods, and results of our study.
-It has been validated through experiments using VGG and ResNet architectures across four datasets, including the ImageNet dataset, demonstrating strong generalization capabilities.
-Guidelines:
-The answer NA means that the abstract and introduction do not include the claims made in the paper.
-The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations.
-A No or NA answer to this question will not be perceived well by the reviewers.
-The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
-It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
-2.
-Limitations
-Question: Does the paper discuss the limitations of the work performed by the authors?
-Answer: [Yes]
-Justification: In the appendix, we have detailed the process of exploring hyperparameters for the SMA.
-The experiments clearly show that the performance of SMA modules is influenced by hyperparameters.
-Furthermore, the AZO regularization method depends on the attention weights of the SMA module.
-Guidelines:
-The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
-The authors are encouraged to create a separate "Limitations" section in their paper.
-The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally).
-The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
-The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs.
-In general, empirical results often depend on implicit assumptions, which should be articulated.
-The authors should reflect on the factors that influence the performance of the approach.
-For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting.
-Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
-The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
-If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
-While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper.
-The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community.
-Reviewers will be specifically instructed to not penalize honesty concerning limitations.
-3.
-Theory Assumptions and Proofs
-Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
-Answer: [N/A]
-Justification: Our work does not belong to the field of deep learning theory research.
-Guidelines:
-The answer NA means that the paper does not include theoretical results.
-All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
-All assumptions should be clearly stated or referenced in the statement of any theorems.
-The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
-Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
-Theorems and Lemmas that the proof relies upon should be properly referenced.
-4.
-Experimental Result Reproducibility
-Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
-Answer: [Yes]
-Justification: In the main text, we have extensively described the complete structure of our proposed method using visual and mathematical representations.
-Additionally, we have listed all hyperparameters for the experiments in the appendix.
-Moreover, we have made our source code publicly available.
-Guidelines:
-The answer NA means that the paper does not include experiments.
-If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
-If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
-Depending on the contribution, reproducibility can be accomplished in various ways.
-For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model.
-In general.
-releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
-While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution.
-For example
-If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
-If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
-If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
-We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility.
-In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
-5.
-Open access to data and code
-Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
-Answer: [Yes]
-Justification: Our experiments were conducted using publicly available datasets. Additionally, we have made all source code and weight files publicly accessible.
-Guidelines:
-The answer NA means that paper does not include experiments requiring code.
-Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
-While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer.
-Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
-The instructions should contain the exact command and environment needed to run to reproduce the results.
-See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
-The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
-The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines.
-If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
-At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
-Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
-6.
-Experimental Setting/Details
-Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
-Answer: [Yes]
-Justification: All hyperparameters are detailed in the appendix.
-Guidelines:
-The answer NA means that the paper does not include experiments.
-The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
-The full details can be provided either with the code, in appendix, or as supplemental material.
-7.
-Experiment Statistical Significance
-Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
-Answer: [Yes]
-Justification: We conducted numerous experiments during the hyperparameter exploration process for the SMA and AZO, recognizing the significance of these for a universal module.
-The results in Tab.
-1 of the main text and Tab.
-8 in the appendix are averages from ten experiments using non-fixed random seeds.
-Similarly, the data shown in Fig.
-5 of the appendix were derived from five experiments with non-fixed random seeds, including error bars to depict variability.Due to constraints on experimental costs and computational resources, we limited our testing to one experiment each on the ImageNet dataset and mainstream neural architecture datasets.
-However, we have carefully preserved the model weights for future reference and use.
-Guidelines:
-The answer NA means that the paper does not include experiments.
-The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
-The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
-The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
-The assumptions made should be given (e.g., Normally distributed errors).
-It should be clear whether the error bar is the standard deviation or the standard error of the mean.
-It is OK to report 1-sigma error bars, but one should state it.
-The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
-For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
-If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
-8.
-Experiments Compute Resources
-Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
-Answer: [Yes]
-Justification: In the main text, we detailed the specific configurations of the three servers used, as well as the service providers for two of the cloud servers.
-All experiments and preliminary work described in this article, including the SMAL module that focuses solely on multiscale interactions without considering spatiotemporal dynamics, have been made open-source.
-Guidelines:
-The answer NA means that the paper does not include experiments.
-The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
-The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
-The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
-9.
-Code Of Ethics
-Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
-Answer: [Yes]
-Justification: Our research, while inspired by the biological brain, does not involve real organisms and contains no generative content, ensuring compliance with ethical standards.
-Guidelines:
-The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
-If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
-The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
-10.
-Broader Impacts
-Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
-Answer: [No]
-Justification: We believe that our proposed method will drive the deployment of SNNs on brain like chips and the study of high-performance SNN models, which we consider to have a social impact.
-However, we did not discuss this in the main text because brain like chips are still in the research stage and their application fields are more inclined towards professional fields.
-We believe that our research has limited social impact beyond the academic field, so we did not discuss it.
-Guidelines:
-The answer NA means that there is no societal impact of the work performed.
-If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
-Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
-The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments.
-However, if there is a direct path to any negative applications, the authors should point it out.
-For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation.
-On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
-The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
-If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
-11.
-Safeguards
-Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
-Answer: [N/A]
-Justification: Our research is purely academic and does not involve generative studies; therefore, it carries no associated risks.
-Guidelines:
-The answer NA means that the paper poses no such risks.
-Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
-Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
-We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
-12.
-Licenses for existing assets
-Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
-Answer: [Yes]
-Justification: The four datasets we use are publicly available for academic research, and we have also cited them.
-Guidelines:
-The answer NA means that the paper does not use existing assets.
-The authors should cite the original paper that produced the code package or dataset.
-The authors should state which version of the asset is used and, if possible, include a URL.
-The name of the license (e.g., CC-BY 4.0) should be included for each asset.
-For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
-If assets are released, the license, copyright information, and terms of use in the package should be provided.
-For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets.
-Their licensing guide can help determine the license of a dataset.
-For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
-If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
-13.
-New Assets
-Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
-Answer: [Yes]
-Justification: Yes, we have made our proposed model publicly available, and all related documentation has also been released.
-Guidelines:
-The answer NA means that the paper does not release new assets.
-Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
-The paper should discuss whether and how consent was obtained from people whose asset is used.
-At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
-14.
-Crowdsourcing and Research with Human Subjects
-Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
-Answer: [N/A]
-Justification: This study did not involve any subjects.
-Guidelines:
-The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
-Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
-According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
-15.
-Institutional Review Board (IRB) Approvals or Equivalent for Research with Human Subjects
-Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
-Answer: [N/A]
-Justification: This study did not involve any subjects.
-Guidelines:
-The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
-Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research.
-If you obtained IRB approval, you should clearly state this in the paper.
-We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
-For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.